In a local restaurant on taco Tuesday, the manager kept track of what the customers ordered. Of the 100 customers, 70 ordered tacos, 50 ordered margaritas, and 30 ordered both tacos and margaritas.

-Find the probability that a customer ordered a margarita, given that s/he ordered a taco.
Answer #1

P(tacos) = 70/100

P(margaritas) = 50/100

P(tacos and margaritas) = 30/100

Using conditional probability formula,

P(Margaritas | Tacos) = P(tacos and margaritas) / P(tacos)

= (30/100) / ( 70 / 100)

= 30/70

= 3/7 ( OR = 0.4286)